OpenStudy (anonymous):

Simplify the following expression. x − 6y + 2z − 5x − 3y + 8z

OpenStudy (lalaly):

10z-4x-9y

OpenStudy (anonymous):

combine like terms. is that clear?

OpenStudy (anonymous):

\[x-6y-2z-5x-3y+8z\] like terms are \[x-5x\] \[-6y-3y\] \[-2z+8z\]

OpenStudy (anonymous):

and \[x-5x=-4x\] \[-6y-3y=-9y\] \[2z+8z=10z\]

OpenStudy (anonymous):

so 'final answer ' is \[-4x-9y+10z\]